OpenStudy (anonymous):

Find the time required for an investment of 5000 dollars to grow to 9000 dollars at an interest rate of 7.5 percent per year, compounded quarterly.

OpenStudy (anonymous):

\[9000=5000\times \left(1+\frac{.075}{4}\right)^{4t}\] solve for \(t\)

OpenStudy (anonymous):

any idea how to do this?

OpenStudy (anonymous):

got it! thanks!

OpenStudy (anonymous):

@satellite73